# Personalized Predictions of Glioblastoma Infiltration: Mathematical Models, Physics-Informed Neural Networks and Multimodal Scans

### v2

@@ 3 Results
    Data from 8 patients (P1-P8), primary tumors and first detected recurrence, from (Lipková et al., 2019). All the patients have T1Gd and FLAIR scans at recurrence.
@@
We demonstrate that using segmentation alone achieves results comparable to those using segmentation and FET, in contrast to the results in (Lipková et al., 2019) where FET was needed to more accurately determine spatial distributions of the cells, particularly near the tumor center. This finding is noteworthy as FET-PET data is not widely available from the clinic, thereby accentuating the practicality of our approach.
@@ 3.1 Synthetic data
The PDEs with the inferred parameters can be used to predict tumor evolution at arbitrary times, before or after the time of imaging. Directly evaluating the PINN at time beyond the range of the temporal domain used in training tends to yield poor results, although this can be remedied by re-sampling collocation points and re-training the network with the PDE loss, or by novel methods to enhance the extrapolation capabilities of PINN (Kim et al., 2020). It should be noted that our method only permits the use of normalized time as the time scale remains unknown. Figure 4 demonstrates the predicted dynamics of tumor cell density $u^{FDM}$ at normalized times $\bar{t}$ = 0.6, 0.8, 1.0, 1.2, with data provided at $\bar{t}=1$. The first row shows the 3D isosurfaces of $u^{FDM}$ at 1% and 30%. The second row shows the predicted contours and those of the GT. The predicted dynamics of tumor growth align closely with the GT because the estimated parameters are close to the GT parameters.
@@ 3.2 Patient tumors
In this section, our method is applied to data from 24 patients: P1-P8 and Q1-Q16. As an example, Figure 5 demonstrates the results of the inference for patient P5 using SEG and FET+SEG as data. The estimated parameters are given in Table 4. The first row presents the preoperative scans (a) T1Gd (b) FLAIR, (c) the pre-processed FET (as described in 2.1.4) and (d) the original FET data. The second and third rows exhibit the results for $u^{FDM}$ using FET+SEG and SEG data, respectively. Upon comparing columns 1, 2, and 4, it is evident that the estimations using FET+SEG and SEG result in similar predicted tumor cell densities, and the predicted segmentations align closely with the actual segmentations. The Radiation Therapy Oncology Group (RTOG) Clinical Target Volume (CTV), denoted as $\rm CTV^{\rm RTOG}$, is defined as the 2 cm margin around the FLAIR segmentation (Gilbert et al. (2013)). The personalized CTV, denoted as $\rm CTV^{\rm P}$, is defined by thresholding the predicted tumor cell density at $u^{FDM}=1\%$ across all patients (blue contours in Figs. 5 (f) and (j)). This threshold was selected so that the efficiency of the $\rm CTV^{\rm P}$ is comparable to that of the $\rm CTV^{\rm RTOG}$ (green contours in Figs. 5 (f) and (j)). Figures 5(f) and (j) reveal that both $\rm CTV^{\rm P}$ and $\rm CTV^{\rm RTOG}$ encompass the visible tumor, although $\rm CTV^{\rm P}$ is smaller in size compared to $\rm CTV^{\rm RTOG}$. In Figure 5(g), the predicted FET signal using the FET+SEG data is displayed, and is highly correlated with the actual FET signal (see D Table 4).
@@
Figure 7 shows the results for patients P1-P8 using SEG data. The estimated model parameters can be found in D Table 4. The T1Gd and FLAIR images are shown in Figs. 7 (A) and (B), and the predictions are shown in Figs. 7 (C-F). All patients received standard treatment, which consists of immediate tumor resection, followed by combined radiotherapy and chemotherapy, and are monitored for recurrence regularly. Fig. 7 (e) shows the post-operative FLAIR scan with segmentations of the first detected tumor recurrences. The postoperative scans are registered to the preoperative T1Gd scans by rigid transformations (Lipková et al., 2019). The figures also plot the $\rm CTV^{\rm P}$ (blue), indicative of the predicted tumor infiltration pathways, and $\rm CTV^{\rm RTOG}$ (green) for comparison. All the $\rm CTV^{\rm P}$ and $\rm CTV^{\rm RTOG}$ cover the recurrent tumor reasonably well. Further, for patient P5, $\rm CTV^{\rm P}$ follows the contour of the recurrent tumor. For patient P8, we see infiltration into the other brain hemisphere beyond the $\rm CTV^{\rm P}$ and $\rm CTV^{\rm RTOG}$. However, our predicted $\rm CTV^{\rm P}$ seems to capture the infiltration pathway. The estimated parameters and the metrics, including DICE scores of the segmentations and correlations with FET signals, corresponding to Fig. 7 SEG are reported in D, Table 4. The results of FET+SEG are reported in D, Fig. 15 and Table 4.
@@
For a quantitative evaluation of the personalized CTVs, we follow Lipková et al. (2019) and use the CTV volume and efficiency of the CTV in capturing tumor recurrence as two criteria. A smaller CTV means that less healthy brain tissue is irradiated. The CTV efficiency is defined as the percentage of the recurrent tumor covered by the CTV:
@@
where $V^{CTV}$ is the CTV region, $V^{REC}$ is the FLAIR segmentation (tumor core and edema region) of the recurrent tumor and ||$\cdot$|| ⋅ | denotes the volume. Figure 8 compares the volumes and efficiencies of RTOG CTV (blue) with the $\rm CTV^{\rm P}$ (1% isosurface of predicted tumor cell density $u^{FDM}$) using different parameter inference results, FET+SEG (red), SEG (yellow), from patient data P1-P8 and Q1-Q3 for which we have both T1Gd and FLAIR images at recurrence. For patients P1-P8, we also include the personalized CTVs based on the maximum a posteriori (MAP) estimates from Lipková et al. (2019) (purple).
The $\rm CTV^{\rm P}$ volumes and efficiencies using FET+SEG and SEG data from our PINN-based inference are very similar and except for patients P2 and P8, where the volumes of the $\rm CTV^{\rm P}$ from Lipková et al. (2019) are significantly smaller, the volumes and efficiencies using our PINN-based approach are similar to those from Lipková et al. (2019), which is much more expensive to compute. Compared to $\rm CTV^{\rm RTOG}$, the $\rm CTV^{\rm P}$ from our study and from Lipková et al. (2019) exhibit reduced or similar volumes while maintaining efficiencies comparable to $\rm CTV^{\rm RTOG}$. Patient Q13 is an outlier, as the recurrent tumor has an exceptionally large edema region, leading to low efficiency for all CTVs. We discuss possible explanations for this in D and Fig. 21.
@@ 3.3 Discussion
We briefly discuss the merits and limitations of our PINN approach, compared with other methods. One of the strengths of Bayesian methods is their ability to quantify uncertainties, while PINNs provide deterministic outputs based on the minimizer of the loss function and do not inherently provide uncertainty estimates, although some methods have been proposed to integrate uncertainty quantification into PINNs (Yang et al., 2021). However, Bayesian methods are computationally expensive as they require solving the forward PDE problem many times to estimate the posterior distribution of the parameters, which might takes up to days (Lipková et al., 2019; Ezhov et al., 2023). Our PINN implementation on TensorFLow (Abadi et al., 2015) takes about 30 min for the pre-training step and 5 min for the fine-tuning step on a single GPU (Nvidia Quadro RTX 8000).
@@
Conventional PDE-constrained optimization methods, such as the adjoint method, are very efficient and solve the PDE using numerical solvers, which come with well-understood error estimators, as well as convergence and stability guarantees (Hogea et al., 2008; Gholami et al., 2016; Subramanian et al., 2020b, a, 2023). In contrast, the training of PINNs remains an active area of research (Karniadakis et al., 2021; Krishnapriyan et al., 2021; Wang et al., 2022b; Hao et al., 2023; Chen and Hao, 2022). One notable advantage of PINNs is their flexibility and simplicity in implementation: modifying the model primarily involves changing the loss function. In contrast, the adjoint method often requires derivation of the adjoint equation, careful regularization, and implementation of the adjoint solver, adding to the technical demands.
@@
Another flexible machine learning approach, known as ODIL (Karnakov et al., 2022), was recently developed to estimate parameters of the Fisher-KPP PDE using MRI and FET-PET scans for GBM (Balcerak et al., 2023). Like the PINN, this method also minimizes a loss function that contains data and the PDE residual. In ODIL the residual of the PDE is calculated using a grid-based discretization of the PDE and the variables to be minimized are the solutions at the grid points in 4D (space+time) rather than weights of a neural network as in the PINN. Although the number of grid points can be very large, the multigrid method and the sparsity of the discrete systems can be used to improve efficiency. While promising, the current implementation uses downsampling of the medical images.
@@
Another class of methods aims to learn the mapping between image and parameters directly, either by training a neural network (Ezhov et al., 2023), or by posing the inverse problem as a database query task (Ezhov et al., 2022b). These methods require a large number of forward PDE solves to generate a synthetic dataset and can take days to compute, which is generally not a concern in practice as it is done offline. The inference is extremely fast since it only requires evaluating the neural network or a for-loop search. In (Ezhov et al., 2023), 100,000 synthetic tumors were generated for training, while the inference was done in about 5 minutes. However, this approach does not allow any patient specific fine-tuning, thus might have larger errors if the patient data is significantly different from the training data.
@@
One might wonder if using an averaged set of the characteristic parameters can be used to pre-train a single network for all patients. We tested this (see H) and found that such a pre-trained single network does not work well with our current framework because the tumors have different scales, as demonstrated by the wide range of patient-specific parameters. Without the patient-specific characteristic parameters, the pre-trained network can be far away from the solution, leading to difficulty in fine-tuning. Nevertheless, the idea of a common pre-trained network for all patients is appealing as it can significantly increase the inference speed. Consequently, in future work we will explore other approaches for developing a single network that can be used for all patients, including transfer learning (Xu et al., 2023; Desai et al., 2021; Gao et al., 2022) and neural operators (Li et al., 2020; Lu et al., 2021a)
@@ 4 Conclusion
The pre-operative MRI scan of patient P1-P8 are available from Lipková et al. (2019). Other data will be made available on request. The PINN code is available at https://github.com/Rayzhangzirui/pinn
@@ Acknowledgments
R.Z.Z and J.S.L thank Jana Lipkova, Peter Chang, and Xiaohui Xie for helpful discussions, and Babak Shahbaba for the GPU resources. J.S.L. acknowledges National Institutes of Health for partial support through grant nos. 1U54CA217378-01A1 for a National Center in Cancer Systems Biology at UC Irvine and P30CA062203 for the Chao Family Comprehensive Cancer Center at UC Irvine. In addition, J.S.L. acknowledges support from DMS-1763272 and the Simons Foundation (594598QN) for an NSF-Simons Center for Multiscale Cell Fate Research, and NSF grants DMS-1936833 and DMS-1953410. BM and MB acknowledge support from the Helmut Horten Foundation.
@@ Appendix A Training of PINN
In Fig. 11, we show an example of the training loss for patient P5 with FET+SEG. The left figure shows the training loss for pre-training with residual loss and $u^{FDM}$. The right figure shows the training loss for fine-tuning with segmentation loss and PET loss. The $\mathcal{L}_{\rm SEG}$ is weighted by 0.001 and $\mathcal{L}_{\rm FET}$ (if necessary) is weighted by 0.01, while all the other losses are weighted by 1. These weights are chosen such that the magnitude of the losses are comparable, and are constant throughout the training. While we experimented with some adaptive weighting schemes, which balance different loss terms according to coefficients of variation (Groenendijk et al., 2020), by the norm of the gradients (Chen et al., 2018; Wang et al., 2021), or the standard deviation of the gradients (Maddu et al., 2022), using the fixed weights listed above achieved the best results across all synthetic and patient datasets, without significantly increasing the training time.
@@ Appendix B Distribution of collocation points
We explain how we sample the collocation points for the residual loss, which affects the training efficiency and accuracy. In Lu et al. (2021b), an adaptive residual refinement technique is used to adaptively increase the density of collocation points in the region where the residual is large. In Wang et al. (2022a), the residuals at early times are adaptively given higher weights. We adopt a similar approach in our work here. We know that the solution of the PDE will have larger gradients at early times, and the solution will be concentrated at the center of the tumor.
The distribution of the collocation points is illustrated in Fig. 12. Let ${\mathbf{x}_{0}}$ be the centroid of $y^{\rm T1Gd}$, choose the radius R such that the spherical ball $B$ with center ${\mathbf{x}_{0}}$ and radius R encloses the region where $\bar{u}^{FDM}>0.001$. Next, we sample the grid points $\mathbf{x}_{i}$ in $C$ such that $\mathbf{x}_{i}$ are dense in the center, sparse near the boundary, and $t_{i}$ are dense at early times and sparse at late times, because the solution $u$ has larger gradients in the tumor center and at early times. More specifically, $t_{i}$ is sampled from a truncated exponential probability density function $\lambda e^{-\lambda t}$, with $\lambda=0.5$ and $t\in[0,1]$. Each spatial coordinate $\mathbf{x}_{i}$ is resampled with weight $1/(r_{i}^{2})$ in 3d, where $r_{i}$ is the distance from $\mathbf{x}_{i}$ to the tumor center ${\mathbf{x}_{0}}$. We experimented with other static distributions of the collocation points, including uniformly sampling the temporal and spatial coordinates, or densely sampling the spatial coordinates at the transition region between white and gray matter (where the diffusion coefficient changes sharply), and an adaptive residual refinement technique following (Lu et al., 2021b), which recruits the 100 testing points with the largest residual loss into the training points every 1000 iterations, but we did not find significant improvement in accuracy.

### v3

@@ 3 Results
    Data from 8 patients (P1-P8), primary tumors and first detected recurrence, from [43]. All the patients have T1Gd and FLAIR scans at recurrence.
@@
We demonstrate that using segmentation alone achieves results comparable to those using segmentation and FET, in contrast to the results in [43] where FET was needed to more accurately determine spatial distributions of the cells, particularly near the tumor center. This finding is noteworthy as FET-PET data is not widely available from the clinic, thereby accentuating the practicality of our approach.
@@ 3.1 Synthetic data
The PDEs with the inferred parameters can be used to predict tumor evolution at arbitrary times, before or after the time of imaging. Directly evaluating the PINN at time beyond the range of the temporal domain used in training tends to yield poor results, although this can be remedied by re-sampling collocation points and re-training the network with the PDE loss, or by novel methods to enhance the extrapolation capabilities of PINN [34]. It should be noted that our method only permits the use of normalized time as the time scale remains unknown. Figure 4 demonstrates the predicted dynamics of tumor cell density $u^{FDM}$ at normalized times $\bar{t}$ = 0.6, 0.8, 1.0, 1.2, with data provided at $\bar{t}=1$. The first row shows the 3D isosurfaces of $u^{FDM}$ at 1% and 30%. The second row shows the predicted contours and those of the GT. The predicted dynamics of tumor growth align closely with the GT because the estimated parameters are close to the GT parameters.
@@ 3.2 Patient tumors
In this section, our method is applied to data from 24 patients: P1-P8 and Q1-Q16. As an example, Figure 5 demonstrates the results of the inference for patient P5 using SEG and FET+SEG as data. The estimated parameters are given in Table 4. The first row presents the preoperative scans (a) T1Gd (b) FLAIR, (c) the pre-processed FET (as described in 2.1.4) and (d) the original FET data. The second and third rows exhibit the results for $u^{FDM}$ using FET+SEG and SEG data, respectively. Upon comparing columns 1, 2, and 4, it is evident that the estimations using FET+SEG and SEG result in similar predicted tumor cell densities, and the predicted segmentations align closely with the actual segmentations. The Radiation Therapy Oncology Group (RTOG) Clinical Target Volume (CTV), denoted as $\rm CTV^{\rm RTOG}$, is defined as the 2 cm margin around the FLAIR segmentation (Gilbert et al. [20]). The personalized CTV, denoted as $\rm CTV^{\rm P}$, is defined by thresholding the predicted tumor cell density at $u^{FDM}=1\%$ across all patients (blue contours in Figs. 5 (f) and (j)). This threshold was selected so that the efficiency of the $\rm CTV^{\rm P}$ is comparable to that of the $\rm CTV^{\rm RTOG}$ (green contours in Figs. 5 (f) and (j)). Figures 5(f) and (j) reveal that both $\rm CTV^{\rm P}$ and $\rm CTV^{\rm RTOG}$ encompass the visible tumor, although $\rm CTV^{\rm P}$ is smaller in size compared to $\rm CTV^{\rm RTOG}$. In Figure 5(g), the predicted FET signal using the FET+SEG data is displayed, and is highly correlated with the actual FET signal (see D Table 4).
@@
Figure 7 shows the results for patients P1-P8 using SEG data. The estimated model parameters can be found in D Table 4. The T1Gd and FLAIR images are shown in Figs. 7 (A) and (B), and the predictions are shown in Figs. 7 (C-F). All patients received standard treatment, which consists of immediate tumor resection, followed by combined radiotherapy and chemotherapy, and are monitored for recurrence regularly. Fig. 7 (e) shows the post-operative FLAIR scan with segmentations of the first detected tumor recurrences. The postoperative scans are registered to the preoperative T1Gd scans by rigid transformations [43]. The figures also plot the $\rm CTV^{\rm P}$ (blue), indicative of the predicted tumor infiltration pathways, and $\rm CTV^{\rm RTOG}$ (green) for comparison. All the $\rm CTV^{\rm P}$ and $\rm CTV^{\rm RTOG}$ cover the recurrent tumor reasonably well. Further, for patient P5, $\rm CTV^{\rm P}$ follows the contour of the recurrent tumor. For patient P8, we see infiltration into the other brain hemisphere beyond the $\rm CTV^{\rm P}$ and $\rm CTV^{\rm RTOG}$. However, our predicted $\rm CTV^{\rm P}$ seems to capture the infiltration pathway. The estimated parameters and the metrics, including DICE scores of the segmentations and correlations with FET signals, corresponding to Fig. 7 SEG are reported in D, Table 4. The results of FET+SEG are reported in D, Fig. 15 and Table 4.
@@
For a quantitative evaluation of the personalized CTVs, we follow Lipková et al. [43] and use the CTV volume and efficiency of the CTV in capturing tumor recurrence as two criteria. A smaller CTV means that less healthy brain tissue is irradiated. The CTV efficiency is defined as the percentage of the recurrent tumor covered by the CTV:
@@
where $V^{CTV}$ is the CTV region, $V^{REC}$ is the FLAIR segmentation (tumor core and edema region) of the recurrent tumor and ||$\cdot$|| ⋅ | denotes the volume. Figure 8 compares the volumes and efficiencies of RTOG CTV (blue) with the $\rm CTV^{\rm P}$ (1% isosurface of predicted tumor cell density $u^{FDM}$) using different parameter inference results, FET+SEG (red), SEG (yellow), from patient data P1-P8 and Q1-Q3 for which we have both T1Gd and FLAIR images at recurrence. For patients P1-P8, we also include the personalized CTVs based on the maximum a posteriori (MAP) estimates from Lipková et al. [43] (purple).
The $\rm CTV^{\rm P}$ volumes and efficiencies using FET+SEG and SEG data from our PINN-based inference are very similar and except for patients P2 and P8, where the volumes of the $\rm CTV^{\rm P}$ from Lipková et al. [43] are significantly smaller, the volumes and efficiencies using our PINN-based approach are similar to those from Lipková et al. [43], which is much more expensive to compute. Compared to $\rm CTV^{\rm RTOG}$, the $\rm CTV^{\rm P}$ from our study and from Lipková et al. [43] exhibit reduced or similar volumes while maintaining efficiencies comparable to $\rm CTV^{\rm RTOG}$. Patient Q13 is an outlier, as the recurrent tumor has an exceptionally large edema region, leading to low efficiency for all CTVs. We discuss possible explanations for this in D and Fig. 21.
@@ 3.3 Discussion
We briefly discuss the merits and limitations of our PINN approach, compared with other methods. One of the strengths of Bayesian methods is their ability to quantify uncertainties, while PINNs provide deterministic outputs based on the minimizer of the loss function and do not inherently provide uncertainty estimates, although some methods have been proposed to integrate uncertainty quantification into PINNs [87]. However, Bayesian methods are computationally expensive as they require solving the forward PDE problem many times to estimate the posterior distribution of the parameters, which might takes up to days [43, 15]. Our PINN implementation on TensorFLow [1] takes about 30 min for the pre-training step and 5 min for the fine-tuning step on a single GPU (Nvidia Quadro RTX 8000).
@@
Conventional PDE-constrained optimization methods, such as the adjoint method, are very efficient and solve the PDE using numerical solvers, which come with well-understood error estimators, as well as convergence and stability guarantees [26, 19, 74, 73, 71]. In contrast, the training of PINNs remains an active area of research [33, 37, 83, 23, 7]. One notable advantage of PINNs is their flexibility and simplicity in implementation: modifying the model primarily involves changing the loss function. In contrast, the adjoint method often requires derivation of the adjoint equation, careful regularization, and implementation of the adjoint solver, adding to the technical demands.
@@
Another flexible machine learning approach, known as ODIL [32], was recently developed to estimate parameters of the Fisher-KPP PDE using MRI and FET-PET scans for GBM [3]. Like the PINN, this method also minimizes a loss function that contains data and the PDE residual. In ODIL the residual of the PDE is calculated using a grid-based discretization of the PDE and the variables to be minimized are the solutions at the grid points in 4D (space+time) rather than weights of a neural network as in the PINN. Although the number of grid points can be very large, the multigrid method and the sparsity of the discrete systems can be used to improve efficiency. While promising, the current implementation uses downsampling of the medical images.
@@
Another class of methods aims to learn the mapping between image and parameters directly, either by training a neural network [15], or by posing the inverse problem as a database query task [14]. These methods require a large number of forward PDE solves to generate a synthetic dataset and can take days to compute, which is generally not a concern in practice as it is done offline. The inference is extremely fast since it only requires evaluating the neural network or a for-loop search. In [15], 100,000 synthetic tumors were generated for training, while the inference was done in about 5 minutes. However, this approach does not allow any patient specific fine-tuning, thus might have larger errors if the patient data is significantly different from the training data.
@@
One might wonder if using an averaged set of the characteristic parameters can be used to pre-train a single network for all patients. We tested this (see H) and found that such a pre-trained single network does not work well with our current framework because the tumors have different scales, as demonstrated by the wide range of patient-specific parameters. Without the patient-specific characteristic parameters, the pre-trained network can be far away from the solution, leading to difficulty in fine-tuning. Nevertheless, the idea of a common pre-trained network for all patients is appealing as it can significantly increase the inference speed. Consequently, in future work we will explore other approaches for developing a single network that can be used for all patients, including transfer learning [85, 11, 18] and neural operators [42, 46]
@@ 4 Conclusion
The pre-operative MRI scan of patient P1-P8 are available from Lipková et al. [43]. Other data will be made available on request. The PINN code is available at https://github.com/Rayzhangzirui/pinn
@@ Acknowledgments
R.Z.Z and J.S.L thank Jana Lipkova, Peter Chang, and Xiaohui Xie for helpful discussions, and Babak Shahbaba for the GPU resources. J.S.L. acknowledges National Institutes of Health for partial support through grant nos. 1U54CA217378-01A1 for a National Center in Cancer Systems Biology at UC Irvine and P30CA062203 for the Chao Family Comprehensive Cancer Center at UC Irvine. In addition, J.S.L. acknowledges support from DMS-1763272, DMS-1936833, DMS-1953410, DMS-2309800 and the Simons Foundation (594598QN) for an NSF-Simons Center for Multiscale Cell Fate Research. BM and MB acknowledge support from the Helmut Horten Foundation.
@@ Appendix A Training of PINN
In Fig. 11, we show an example of the training loss for patient P5 with FET+SEG. The left figure shows the training loss for pre-training with residual loss and $u^{FDM}$. The right figure shows the training loss for fine-tuning with segmentation loss and PET loss. The $\mathcal{L}_{\rm SEG}$ is weighted by 0.001 and $\mathcal{L}_{\rm FET}$ (if necessary) is weighted by 0.01, while all the other losses are weighted by 1. These weights are chosen such that the magnitude of the losses are comparable, and are constant throughout the training. While we experimented with some adaptive weighting schemes, which balance different loss terms according to coefficients of variation [21], by the norm of the gradients [9, 82], or the standard deviation of the gradients [51], using the fixed weights listed above achieved the best results across all synthetic and patient datasets, without significantly increasing the training time.
@@ Appendix B Distribution of collocation points
We explain how we sample the collocation points for the residual loss, which affects the training efficiency and accuracy. In Lu et al. [47], an adaptive residual refinement technique is used to adaptively increase the density of collocation points in the region where the residual is large. In Wang et al. [81], the residuals at early times are adaptively given higher weights. We adopt a similar approach in our work here. We know that the solution of the PDE will have larger gradients at early times, and the solution will be concentrated at the center of the tumor.
The distribution of the collocation points is illustrated in Fig. 12. Let ${\mathbf{x}_{0}}$ be the centroid of $y^{\rm T1Gd}$, choose the radius R such that the spherical ball $B$ with center ${\mathbf{x}_{0}}$ and radius R encloses the region where $\bar{u}^{FDM}>0.001$. Next, we sample the grid points $\mathbf{x}_{i}$ in $C$ such that $\mathbf{x}_{i}$ are dense in the center, sparse near the boundary, and $t_{i}$ are dense at early times and sparse at late times, because the solution $u$ has larger gradients in the tumor center and at early times. More specifically, $t_{i}$ is sampled from a truncated exponential probability density function $\lambda e^{-\lambda t}$, with $\lambda=0.5$ and $t\in[0,1]$. Each spatial coordinate $\mathbf{x}_{i}$ is resampled with weight $1/(r_{i}^{2})$ in 3d, where $r_{i}$ is the distance from $\mathbf{x}_{i}$ to the tumor center ${\mathbf{x}_{0}}$. We experimented with other static distributions of the collocation points, including uniformly sampling the temporal and spatial coordinates, or densely sampling the spatial coordinates at the transition region between white and gray matter (where the diffusion coefficient changes sharply), and an adaptive residual refinement technique following [47], which recruits the 100 testing points with the largest residual loss into the training points every 1000 iterations, but we did not find significant improvement in accuracy.